OpenStudy (marsmcnugget):

What is the value of the expression? 40 ÷ [24 – 4 • (2 + 3)]

OpenStudy (michele_laino):

first step: we have to compute 2+3=...?

OpenStudy (marsmcnugget):

5

OpenStudy (michele_laino):

right! second step: we have to compute 4*5=...?

OpenStudy (michele_laino):

4*(2+3)= 4*5=...?

OpenStudy (marsmcnugget):

20??

OpenStudy (marsmcnugget):

@Michele_Laino

OpenStudy (michele_laino):

ok! now we have to compute this: 24 - 4*(2+3)= 24 - 20=...?

OpenStudy (marsmcnugget):

4?

OpenStudy (michele_laino):

correct! Finally we have to compute this: 40 : )24 - 4*(2+3)) = 40 : 4=...?

OpenStudy (michele_laino):

oops.. 40 : (24 - 4*(2+3)) = 40 : 4=...?

OpenStudy (marsmcnugget):

10

OpenStudy (michele_laino):

that's right! that is your answer!

OpenStudy (marsmcnugget):

thank you :)

OpenStudy (michele_laino):

:)